OpenStudy (fools101):

Hi Guys! Help me !! please! Zeros & Remainder Theorem 1. What is the remainder when 2x2 + 6x + 3 is divided by (x + 3)? A) 39 B) 3 C) 1 D) 0

jimthompson5910 (jim_thompson5910):

Using the remainder theorem, if you divide f(x) by x-r, then that's the same as computing f(r) In this case, r = -3 So f(r) = f(-3) = 2(-3)^2+6(-3)+3 = 2(9)+6(-3)+3 = 18-18+3 = 3 This means that f(-3) = 3 So the remainder of dividing 2x2 + 6x + 3 by (x + 3) is 3 So the answer is choice B

OpenStudy (fools101):

AWWW kk thank u !!!JIM ^ ^

jimthompson5910 (jim_thompson5910):

np